Can some one show me how to solve this problem?
Earth, Jupiter, Saturn, and Uranus all revolve around the sun. Earth takes 1 year, Jupiter 12 years, Saturn 30 years, and Uranus 84 years. How frequently do these four Planets line up with each other?

Answer by stanbon(75887)

Earth, Jupiter, Saturn, and Uranus all revolve around the sun. Earth takes 1 year, Jupiter 12 years, Saturn 30 years, and Uranus 84 years. How frequently do these four Planets line up with each other?
---------------
You need to find the least common multiple of 1,12,30,84
Express each number as a product of primes:
1 = 1
12= 2^2*3
30=2*3*5
84=2^2*3*7
----------------
lcm = 2^2*3*5*7=420
-------------
They will line up once in every 420 years.
==============
Cheers,
Stan H.